# Late Jets, Early Sparks: Illuminating the Pre-Maximum Bumps in Superluminous Supernovae

### v1

@@ 4 Discussion
Cocoon emission following the birth of a compact object has been invoked in several energetic transients, including collapsars (e.g., Nakar 2015; Nakar & Piran 2017) and neutron star mergers (e.g., Lazzati et al. 2017; Gottlieb et al. 2018a, b). However, in general, the cocoon emission in these systems is radiatively inefficient, such that a significant fraction of the deposited jet energy goes into adiabatic losses rather than observable radiation. By contrast, the longer duration of the engines at work in SLSNe, which inject the bulk of their energy over days to weeks instead of seconds to minutes, and correspondingly much lower densities of the supernova ejecta at these late times, reduce the impact of adiabatic losses. In particular, the trapping radius moves inside the JCI on a timescale comparable to the engine duration. This enables an order-unity fraction of the jet luminosity to emerge as radiation.
@@
Such a high radiative efficiency is important because the luminosities of the early bumps seen in SLSN light curves $\sim 10^{43}-10^{45}$ erg s-1 correspond to a sizable fraction $\sim 1-100\%$ of the total energy budget of the engine (Fig. 1). In the case of a magnetar engine, the fraction of the magnetar spin-down energy placed into a jet depends on several uncertain conditions, such as the inclination angle of the magnetic dipole (Margalit et al. 2018, which determines the fraction of the striped pulsar wind that undergoes magnetic reconnection at the wind termination shock; e.g., Lyubarsky 2003; Komissarov 2013) and the efficiency with which any remaining large-scale field avoids instabilities and associated reconnection (Bucciantini et al., 2009; Porth et al., 2013; Mösta et al., 2015). In the case of a black hole engine, the power of the relativistic jet depends on the magnetic field and spin of the black hole (Blandford & Znajek, 1977), while accretion disk outflows can release their energy more isotropically and contribute to powering the bulk of the SLSN emission (e.g., Dexter & Kasen 2013). If the jet engine operates for $\,{t_{\rm eng}}\lesssim t_{b}$, then the jet will be choked deep inside the ejecta, resulting in no powerful breakout and hence absent or weaker associated UV/optical emission. A substantial failed jet fraction is compatible with early maxima being detected in only 3/14 SLSNe-I observed by DES (Angus et al., 2019).
@@
For several days after the jet emerges from the star, the cocoon emission is hot $T_{\rm c}\approx(1-2)\times 10^{4}$ K and peaks in the optical/NUV at an absolute magnitude $M_{\rm AV}\approx-21.5$. Although multi-band measurements of early SLSN bumps which constrain their temperatures are sparse, such data on PS1-10pm indeed indicated high $\approx(2-3)\times 10^{4}$ K temperatures during its early bump phase (Nicholl & Smartt, 2016). Similar temperature evolution was observed from the early maximum in DES14x3taz, which cooled from $\approx 2.5\times 10^{4}$ K at discovery to $\lesssim 10^{4}$ K over roughly a week (Smith et al., 2016).
@@
For a source with $M_{\rm AB}\approx-21$ lasting a few days, we estimate a ULTRASAT horizon distance of $D_{\rm lim}\approx 4$ Gpc. We assume the early NUV emission to be detectable over $\theta_{\rm obs}\approx 0.5\,{\rm rad}$. Given the volumetric rate of SLSNe-I of $\mathcal{R}\approx 40$ Gpc-3 yr-1 (Quimby et al., 2013; Frohmaier et al., 2021; Zhao et al., 2021) at $z\sim 0.2-1$, and assuming that all SLSNe produce early jet breakout emission similar to that we have predicted, the number of detectable sources per year can be estimated as $N_{\rm det}=(2\pi/3)\theta_{\rm obs}^{2}D_{\rm lim}^{3}\mathcal{R}f_{\Omega}\sim 10^{2}\,{\rm yr}^{-1}$. The detection rates will be lower if jets are less powerful than our baseline model, or only successful in a fraction of explosions, as discussed above and supported by SLSNe-I observations (Angus et al., 2019). More direct rate estimates could be made based on a large uniform sample of pre-maximum SLSNe light curve observations (Nicholl & Smartt, 2016). Given its greater depth and similar cadence, many of the ULTRASAT sources will also have optical light curve data from the Vera C. Rubin Observatory (Ivezić et al., 2019), enabling better constraints on the temperature evolution.
@@
An early photometric detection could also enable prompt spectra to be obtained. We predict the earliest phases of the emission could reveal Doppler-broadened velocities extending to $v/c\gtrsim 0.1$ (lower right-hand panel of Fig. 4), at least for observers relatively close to the jet axis. These and other jet breakout signatures, such as luminous X-ray emission for face-on observers (possibly already observed in one SLSN; Levan et al. 2013) or radio afterglow emission from the jet or its cocoon (Coppejans et al. 2018; Eftekhari et al. 2021; Fig. 5), would provide the ultimate “smoking gun” of an engine-powered origin for SLSNe. Our model for powering the early bumps in SLSNe bears some resemblance to the proposal by Kasen et al. (2016), insofar that we are also invoking shock heating of the supernova ejecta due to energy input from a central engine which is substantially delayed after the explosion. However, while Kasen et al. (2016) consider a spherical (or quasi-spherical) shock breakout driven by the high pressure of a nebula (“bubble”) inflated by a central engine (see also Suzuki & Maeda 2016), we have considered the breakout of a tightly collimated jet, closer to the proposal of Margalit et al. (2018). Furthermore, for the nebula-driven shock to reach the ejecta surface, the energy of the central engine must exceed the original kinetic energy of the explosion (Kasen et al., 2016), while collimated jet breakout from homologous ejecta places a less stringent requirement on the engine energy (e.g., Gottlieb & Nakar 2022). The successful breakout of a relativistic jet also comes with additional consequences, such as a synchrotron afterglow from relativistic ejecta, that in general will not be present in the spherical breakout case (however, see Arons 2003; Bucciantini et al. 2011). The outer layers of the ejecta accelerated by a spherical breakout can produce their own synchrotron emission (e.g., Maeda et al. 2023), but exhibit a different light curve and spectral evolution than a jet afterglow, particularly for observers close to the jet axis.
@@
The direct detection of internal high-energy radiation from the jet viewed on-axis would provide an even more dramatic confirmation of the model. The mechanism, radiative efficiency, and spectrum of any internal jet emission are highly uncertain. Nevertheless, we focus on the soft X-ray band, motivated by the X-ray detection of SCP 06F6 (Levan et al., 2013) and the recent launch of the satellite mission Einstein Probe, which reaches a sensitivity in the soft X-ray band across a large fraction of the sky of $F_{\rm lim}\approx 10^{-12}$ erg cm-2 s-1 for a $10^{5}$ s integration (Yuan et al., 2022), comparable to the expected duration of SLSNe jets. Assuming that a fraction $f_{\rm X}$ of the jet power $L_{\rm j}$ is placed into soft X-rays, the resulting detection horizon is $D_{\rm lim}\approx(f_{\rm X}L_{\rm j}/4\pi F_{\rm lim})^{1/2}$. Adopting a beaming fraction $f_{\rm b}\sim 10^{-2}$ similar to those of GRB jets, the number of detectable on-axis SLSNe jets per year can be estimated as:
@@
It is thus possible that Einstein Probe or successor missions could detect on-axis SLSNe-I jets over the coming years, though given the many uncertainties, a lack of such detections is not necessarily constraining.
We thank Sebastian Gomez and Matt Nicholl for sharing data from their recent compilation of SLSNe light curves, and for helpful comments on the manuscript. OG is supported by the Flatiron Research Fellowship. BDM was supported in part by the National Science Foundation (grant No. AST-2009255) and by the NASA Fermi Guest Investigator Program (grant No. 80NSSC22K1574). The Flatiron Institute is supported by the Simons Foundation.

### v2

@@ 4 Discussion
Cocoon emission following the birth of a compact object has been invoked in several energetic transients, including collapsars (e.g., Nakar 2015; Nakar & Piran 2017) and neutron star mergers (e.g., Lazzati et al. 2017; Gottlieb et al. 2018a, b). However, in general, the cocoon emission in these systems is radiatively inefficient, such that a significant fraction of the deposited jet energy goes into adiabatic losses rather than observable radiation. By contrast, the longer duration of the engines at work in SLSNe, which inject the bulk of their energy over days to weeks instead of seconds to minutes, and the correspondingly much lower densities of the SN ejecta at these late times, reduce the impact of adiabatic losses. In particular, the trapping radius moves inside the JCI on a timescale comparable to the engine duration. This enables an order-unity fraction of the jet luminosity to emerge as radiation.
@@
Such a high radiative efficiency is important because the luminosities of the early bumps seen in SLSN light curves $\sim 10^{43}-10^{45}$ erg s-1 correspond to a sizable fraction $\sim 1-100\%$ of the total energy budget of the engine (Fig. 1). In the case of a magnetar engine, the fraction of the magnetar spin-down energy placed into a jet depends on several uncertain conditions, such as the inclination angle of the magnetic dipole (Margalit et al. 2018, which determines the fraction of the striped pulsar wind that undergoes magnetic reconnection at the wind termination shock; e.g., Lyubarsky 2003; Komissarov 2013) and the efficiency with which any remaining large-scale field avoids instabilities and associated reconnection (Bucciantini et al., 2009; Porth et al., 2013; Mösta et al., 2015). In the case of a black hole engine, the power of the relativistic jet depends on the magnetic field and spin of the black hole (Blandford & Znajek, 1977), while accretion disk outflows can release their energy more isotropically and contribute to powering the bulk of the SLSN emission (e.g., Dexter & Kasen 2013). If the jet engine operates for $\,{t_{\rm eng}}\lesssim t_{b}$, then the jet will be choked deep inside the ejecta, resulting in no powerful breakout and hence absent or weaker associated UV/optical emission. A substantial failed jet fraction is compatible with early maxima being detected in only 3/14 Type I SLSNe observed by DES (Angus et al., 2019).
@@
For several days after the jet emerges from the star, the cocoon emission is hot, $T_{\rm c}\gtrsim 3\times 10^{4}$ K, and peaks in the ultraviolet at an absolute magnitude $M_{\rm AB}\approx-20$. Although multi-band measurements of early SLSN bumps which constrain their temperatures are sparse, such data on PS1-10pm indeed indicated high $\approx(2-3)\times 10^{4}$ K temperatures during its early bump phase (Nicholl & Smartt, 2016). Similar temperature evolution was observed from the early maximum in DES14x3taz, which cooled from $\approx 2.5\times 10^{4}$ K at discovery to $\lesssim 10^{4}$ K over roughly a week (Smith et al., 2016).
@@
For a source with $M_{\rm AB}\approx-20$ lasting several days, we estimate an ULTRASAT horizon distance of $D_{\rm lim}\approx 3$ Gpc. We assume the early NUV emission to be detectable over $\theta_{\rm obs}\approx 0.5\,{\rm rad}$. Given the volumetric rate of Type I SLSNe of $\mathcal{R}\approx 40$ Gpc-3 yr-1 (Quimby et al., 2013; Frohmaier et al., 2021; Zhao et al., 2021) at $z\sim 0.2-1$, and assuming that all SLSNe produce early jet breakout emission similar to what we have predicted, the number of detectable sources per year can be estimated as $N_{\rm det}=(2\pi/3)\theta_{\rm obs}^{2}D_{\rm lim}^{3}\mathcal{R}f_{\Omega}\sim 10^{2}\,{\rm yr}^{-1}$. The detection rates will be lower if jets are less powerful than our baseline model, or only successful in a fraction of explosions, as discussed above and supported by Type I SLSNe observations (Angus et al., 2019). More direct rate estimates could be made based on a large uniform sample of premaximum SLSN light curve observations (Nicholl & Smartt, 2016). Given its greater depth and similar cadence, many of the ULTRASAT sources will also have optical light curve data from the Vera C. Rubin Observatory (Ivezić et al., 2019), enabling better constraints on the temperature evolution.
@@
An early photometric detection could also enable prompt spectra to be obtained. We predict that the earliest phases of the emission could reveal Doppler-broadened velocities extending to $v/c\gtrsim 0.1$ (bottom right panel of Fig. 4), at least for observers relatively close to the jet axis. These and other jet breakout signatures, such as luminous X-ray emission for face-on observers (possibly already observed in one SLSN; Levan et al. 2013) or radio afterglow emission from the jet or its cocoon (Coppejans et al. 2018; Eftekhari et al. 2021; Fig. 5), would provide the ultimate “smoking gun” of an engine-powered origin for SLSNe. Our model for powering the early bumps in SLSNe bears some resemblance to the proposal by Kasen et al. (2016), insofar that we are also invoking shock heating of the SN ejecta due to energy input from a central engine which is substantially delayed after the explosion. However, while Kasen et al. (2016) consider emission from a spherical (or quasi-spherical) shock breakout driven by the high pressure of a nebula (“bubble”) inflated by a central engine (see also Suzuki & Maeda 2016), we have considered emission from the interaction of a tightly collimated jet with its cocoon, closer to the proposal of Margalit et al. (2018). Furthermore, for the nebula-driven shock to reach the ejecta surface, the energy of the central engine must exceed the original kinetic energy of the explosion (Kasen et al., 2016), while collimated jet breakout from homologous ejecta places a less stringent requirement on the engine energy (e.g., Gottlieb & Nakar 2022). The successful breakout of a relativistic jet also comes with additional consequences, such as a synchrotron afterglow from relativistic ejecta, that in general will not be present in the spherical breakout case (however, see Arons 2003; Bucciantini et al. 2012). The outer layers of the ejecta accelerated by a spherical breakout can produce their own synchrotron emission (e.g., Maeda et al. 2023), but exhibit a different light curve and spectral evolution than a jet afterglow, particularly for observers close to the jet axis.
@@
The direct detection of internal high-energy radiation from the jet viewed on-axis would provide an even more dramatic confirmation of the model. The mechanism, radiative efficiency, and spectrum of any internal jet emission are highly uncertain. Nevertheless, we focus on the soft X-ray band, motivated by the X-ray detection of SCP 06F6 (Levan et al., 2013) and the recent launch of the satellite mission Einstein Probe, which reaches a sensitivity in the soft X-ray band across a large fraction of the sky of $F_{\rm lim}\approx 10^{-12}$ erg cm-2 s-1 for a $10^{5}$ s integration (Yuan et al., 2022), comparable to the expected duration of SLSN jets. Assuming that a fraction $f_{\rm X}$ of the jet power $L_{\rm j}$ is placed into soft X-rays, the resulting detection horizon is $D_{\rm lim}\approx(f_{\rm X}L_{\rm j}/4\pi F_{\rm lim})^{1/2}$. Adopting a beaming fraction $f_{\rm b}\sim 10^{-2}$ similar to those of GRB jets, the number of detectable on-axis SLSN jets per year can be estimated as:
@@
It is thus possible that Einstein Probe or successor missions could detect on-axis Type I SLSN jets over the coming years, although given the many uncertainties, a lack of such detections is not necessarily constraining.
We thank Sebastian Gomez and Matt Nicholl for sharing data from their recent compilation of SLSN light curves and for helpful feedback on the manuscript. We also thank the referee, Ben Margalit, Raffaella Margutti, and Dan Kasen for useful comments. OG is supported by the Flatiron Research Fellowship. BDM was supported in part by the National Science Foundation (grant No. AST-2009255) and by the NASA Fermi Guest Investigator Program (grant No. 80NSSC22K1574). The Flatiron Institute is supported by the Simons Foundation. The computations in this work were, in part, run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation.